Staff:

Lbto:

Arcetri:

Summary.

Tasks for today. This is time dependant and how much time is available today. There are other activities; luci2 mask exchange and pepsi work.
  • Check re-rotator position.
  • Close loop pause/resume version.
  • TBD. Luci images.
  • TBF Bin2 pupil check.

Details.

11:40am

telescope release for SoulSX

Time until 2PM.

Set flat->638

TEST PRESET.

Thu Nov 18 18:45:57.885 2021 aos.PresetAO{FLAO}.started - PresetAO{FLAO}[330] (aomode:ACE-AO,focst:bentGregorianFront,instr:LUCIFER,wfs:FLAOWFS,sox:0[0],soy:0,rox:0[0],roy:0,mag:11,cindex:0,,elev:nan,rotang:nan,grvang:nan,r0:0,skybr:nan,windsp:nan,winddir:nan)
Thu Nov 18 18:45:57.903 2021 aos.info - WFS source set to: FLAOWFS

LOADED THE PUPILS FILE IN THE AO TABLE.

B1:20211117_180043

PUPIL REROTATOR AT 0

OPTIMIZE GAIN.

Server connection up
2021-11-18 18:51:53.867110| -13006 - [AOException] OptimizeGain: FATAL - Cannot detect modal basis loaded on AdSec (-3006) VAR_NOT_FOUND_ERROR

maual gain.

Not oberved the fringes.

TN: 20211118_185449

tn: 190929

12:00 185449

Check the autogain. It is not working proerly

Brandon is looking at the autogain.py code.


2021-11-18 19:22:39.684544| -13006 - [AOException] OptimizeGain: FATAL - invalid syntax (autogain.py, line 21)

Another try.

2021-11-18 19:27:06.609215| -13006 - [AOException] OptimizeGain: FATAL - Cannot detect modal basis loaded on AdSec: M2C@M_ADSEC (-3006) VAR_NOT_FOUND_ERROR

2021-11-18 19:29:06.289151| -13006 - [AOException] OptimizeGain: FATAL - Cannot detect modal basis loaded on AdSec: M2C (-3006) VAR_NOT_FOUND_ERROR

close loop manual gain

TN:20211118_193053. + luci images.

The optimize gain still reporting an issue.

2021-11-18 19:32:55.468489| -13006 - [AOException] OptimizeGain: FATAL - Cannot detect modal basis loaded on AdSec: M2C (-3006) VAR_NOT_FOUND_ERROR

cancel preset.

Thu Nov 18 19:35:48.724 2021 aos.PresetAO{FLAO}.started - PresetAO{FLAO}[332] (aomode:ACE-AO,focst:bentGregorianFront,instr:LUCIFER,wfs:FLAOWFS,sox:0[0],soy:0,rox:0[0],roy:0,mag:11,cindex:0,,elev:nan,rotang:nan,grvang:nan,r0:0,skybr:nan,windsp:nan,winddir:nan)
Thu Nov 18 19:35:48.741 2021 aos.info - WFS source set to: FLAOWFS

2021-11-18 19:41:01.216567| -13006 - [AOException] OptimizeGain: FATAL - Cannot detect modal basis loaded on AdSec: M2C (-3006) VAR_NOT_FOUND_ERROR

autogain from terminal. it didn;t work.

Doug connected discussing this problems...

hexapod at the liit. Homing the hexapod.

1:41PM

ADSC_STOP/START

Thu Nov 18 20:45:11.686 2021 aos.PresetAO{FLAO}.started - PresetAO{FLAO}[335] (aomode:ACE-AO,focst:bentGregorianFront,instr:LUCIFER,wfs:FLAOWFS,sox:0[0],soy:0,rox:0[0],roy:0,mag:11,cindex:0,,elev:nan,rotang:nan,grvang:nan,r0:0,skybr:nan,windsp:nan,winddir:nan)
Thu Nov 18 20:45:11.710 2021 aos.info - WFS source set to: FLAOWFS

app to detect if there is a problem with the message deamon... use the brandon check ...

THIS SEEMS. A GOOD ADD.

Load Enrico's table with the EMGain values optimized for 900Hz, the pupils files changed for bin1 and 4.

Thu Nov 18 21:20:53.895 2021 aos.PresetAO{FLAO}.started - PresetAO{FLAO}[337] (aomode:ACE-AO,focst:bentGregorianFront,instr:LUCIFER,wfs:FLAOWFS,sox:0[0],soy:0,rox:0[0],roy:0,mag:11,cindex:0,,elev:nan,rotang:nan,grvang:nan,r0:0,skybr:nan,windsp:nan,winddir:nan)

preset mag 6

Thu Nov 18 21:25:31.206 2021 aos.AcquireRef{FLAO}.complete - AcquireRef{FLAO}[338] - dx:-0.0726947 dy:1.51872 slNull: f1:0.0 f2:1.0 freq:900 mag:0 nBins:1 nModes:501 r0:nan snMode:0 strehl:0 ttMod:0
Thu Nov 18 21:25:58.492 2021 aos.PresetAO{FLAO}.started - PresetAO{FLAO}[339] (aomode:ACE-AO,focst:bentGregorianFront,instr:LUCIFER,wfs:FLAOWFS,sox:0[0],soy:0,rox:0[0],roy:0,mag:6,cindex:0,,elev:nan,rotang:nan,grvang:nan,r0:0,skybr:nan,windsp:nan,winddir:nan)

optimize ok

gopt ok

Stop and reconfigure with the working environment.

lbto@obs4:2 % ssh -XY AOeng@sxadsec.mountain.lbto.org
AOeng@sxadsec.mountain.lbto.org's password:
\\Last login: Thu Nov 18 21:41:08 2021 from 150.135.245.193
ADOPT_ROOT=/home/aoeng/soul/latest/aoroot
ADOPT_SOURCE=/home/aoeng/soul/latest/source
[AOeng@sxadsec source]$
[AOeng@sxadsec source]$

Last login: Thu Nov 18 21:41:10 2021 from 150.135.245.193
-bash: /home/AOeng/.bashrc: line 32: syntax error near unexpected token `fi'
-bash: /home/AOeng/.bashrc: line 32: `fi'
ADOPT_ROOT=/home/AOeng/soul/latest/aoroot
ADOPT_SOURCE=/home/AOeng/soul/latest/source
[AOeng@soul-sxwfs source]$
[AOeng@soul-sxwfs source]$ w_start

setflat->638

Thu Nov 18 21:47:15.512 2021 aos.PresetAO{FLAO}.started - PresetAO{FLAO}[341] (aomode:ACE-AO,focst:bentGregorianFront,instr:LUCIFER,wfs:FLAOWFS,sox:0[0],soy:0,rox:0[0],roy:0,mag:11,cindex:0,,elev:nan,rotang:nan,grvang:nan,r0:0,skybr:nan,windsp:nan,winddir:nan)
Thu Nov 18 21:47:15.525 2021 aos.info - WFS source set to: FLAOWFS

CCD47 doen't open.

Acquisition start... not star found

2021-11-18 22:01:51.180726| Command canceled
2021-11-18 22:03:39.655558| -13006 - [AOException] centerStar: WARNING - RETRY: No stars found in TV frame (-20001)


revert to pause/resume version.

Last login: Thu Nov 18 21:41:14 2021 from 192.168.58.35
ADOPT_ROOT=/home/aoeng/soul/pause_resume_test_20211011/aoroot
ADOPT_SOURCE=/home/aoeng/soul/pause_resume_test_20211011/ao-supervisor
[AOeng@sxadsec ao-supervisor]$
[AOeng@sxadsec ao-supervisor]$
[AOeng@sxadsec ao-supervisor]$
[AOeng@sxadsec ao-supervisor]$ adsc_start
Side: L
Subsystem: ADSEC

setflat->638

Last login: Thu Nov 18 21:43:34 2021 from 150.135.245.193
ADOPT_ROOT=/home/AOeng/pause_resume_test_20211011/aoroot
ADOPT_SOURCE=/home/AOeng/pause_resume_test_20211011/ao-supervisor
[AOeng@soul-sxwfs ao-supervisor]$

Handover.

3:10PM

Data analysis.

Manual gain. no distiurbance

TN: -193053

BIN1_NODIST-MANUALGAIN_20211118.png

Tracknum 20211118_193053
Closed Loop Yes
Is OK? No - Background correction failed - No compatible luci dark found (i.e. same exposure time or filter or dimensions)
AntiDrift OFF
nphotons/sub/fr -1.000000
Magnitude NaN
Binning 1
Frequency [Hz] 900
Modulation 3.000000
FW1 Dichroic 600-1000 nm
FW2 empty
Telescope elevation 0.000000
Wind speed 0.000000
number of skipped frames 0
lambda [um] 1.646000
exptime [s] 25.024212
framerate [Hz] 0.039961
no. frames 1
SR SE 1.074617
LUCI dark  

-- JuanCarlosGuerra - 18 Nov 2021
Topic revision: r2 - 19 Nov 2021, JuanCarlosGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback